OpenStudy (anonymous):

8x+y=6?find y would it be -2?

OpenStudy (anonymous):

No. you will get y in terms of x: \(8x + y =6\)\[\implies y =-8x + 6\]

OpenStudy (anonymous):

So that is the 'slope intercept' form of your line.

OpenStudy (anonymous):

y=-8x+6?

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

I just subtracted 8x from both sides: 8x + y = 6 -8x - 8x =========== 0x + y = 6 - 8x \[\implies y = -8x + 6\]

OpenStudy (anonymous):

OOOH So because you couldnt get a answer you rewrote it?

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

I did get the answer. I just showed you another way to write it.

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

and the slope of the line?

OpenStudy (anonymous):

The slope is the coefficient on the x variable. (The number in front of the x after you have solved for y).

OpenStudy (anonymous):

So in this case, what is the number in front of the x?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

-8 actually.

OpenStudy (anonymous):

So negative 8 is the slope of the line.

OpenStudy (anonymous):

:O TY!